# The Relational Bottleneck as an Inductive Bias for Efficient Abstraction

### v2

@@ Abstract
A central challenge for cognitive science is to explain how abstract concepts are acquired from limited experience. This effort has often been framed in terms of a dichotomy between connectionist and symbolic cognitive models. Here, we highlight a recently emerging line of work that suggests a novel reconciliation of these approaches, by exploiting an inductive bias that we term the relational bottleneck. We review a family of models that employ this approach to induce abstractions in a data-efficient manner, emphasizing their potential as candidate models for the acquisition of abstract concepts in the human mind and brain.
@@
  • Human learners efficiently acquire abstract concepts from limited experience. The effort to explain this capacity has fueled debate between proponents of symbolic and connectionist approaches, and motivated proposals for hybrid neuro-symbolic systems.
  • A recently emerging approach, that we term the ‘relational bottleneck’ principle, suggests a novel way to bridge the gap. By restricting the flow of information from perceptual to downstream reasoning systems to consist only of relations, the relational bottleneck encourages abstraction and the emergence of symbol-like mechanisms in neural networks.
  • We formulate the relational bottleneck principle in information theoretic terms, and review neural network architectures that implement this principle, displaying rapid learning of relational patterns, and systematic generalization of those patterns to novel inputs.
  • The approach may help to explain a diverse set of phenomena, ranging from cognitive development to capacity limits in cognitive function. The approach is also consistent with findings from cognitive neuroscience, and may offer a useful general principle for designing more powerful artificial learning systems.
@@
Human cognition is characterized by a remarkable ability to transcend the specifics of limited experience to entertain highly general, abstract ideas. Understanding how the mind and brain accomplish this has been a central challenge throughout the history of cognitive science, and a major preoccupation of philosophy before that [1, 2, 3, 4]. Of particular importance is the central role played by relations, which enable human reasoners to abstract away from the details of individual entities and identify higher-order patterns across distinct domains [5, 6]. For instance, when presented with the images in Figure 1, one can easily determine that a common relational pattern (ABA) is displayed on both the left and right, despite the involvement of completely different objects. This capacity to think in terms of relations is a major component underlying the human capacity for fluid reasoning [7, 8], and has been proposed as a key factor distinguishing human intelligence from that of other species [9].
@@
A long tradition in both cognitive science and AI [10, 11, 12] holds that this capacity for abstraction depends on processes akin to symbolic programs. A major appeal of this approach lies in the fact that symbols are, by design, abstracted away from the content to which they refer, thus providing a natural account for the flexibility and systematicity of human concepts [13]. More recently, probabilistic program induction (see Glossary) has provided an account of how symbolic concepts might be learned directly from data [14, 15, 16, 17, 18, 19, 20], by formalizing learning as a search for the program that maximizes the likelihood of the observed data. However, while this approach is capable in principle of representing any possible set of concepts [21], the discovery of these concepts using traditional search methods often proves intractable, making it difficult in practice to identify programs with the richness and complexity of human natural concepts (though see Box 2 for discussion of recent approaches).
@@
An alternative approach, connectionism, has for decades explored how the abstractions associated with human cognition might emerge through experience in neural architectures using general-purpose learning algorithms [22, 23, 24, 25]. This endeavor has recently taken on new relevance, as the success of large language models has demonstrated that it is possible, in some cases, for a human-like capacity for abstraction to emerge given sufficient scaling of both architecture and training data [26, 27, 28, 29]. For instance, it has recently been shown that large language models can solve various analogy problems at a level equal to that of college students [30]. However, the ability of these models to perform abstract tasks (e.g., analogy) depends on exposure to a much larger training corpus than individual humans receive in an entire lifetime [31, 32], thus failing to account for the data efficiency of human concept learning.
@@
In this review, we highlight a recently emerging approach that suggests a novel reconciliation of these two traditions. The central feature of this approach is an inductive bias that we refer to as the relational bottleneck: a constraint that biases neural network models to focus on relations between objects rather than the attributes of individual objects (Box 1). This approach satisfies two key desiderata for models of abstract concept acquisition. First, the approach is capable of rapid (i.e., data-efficient) acquisition of abstract relational concepts. Second, the approach does not require access to a set of pre-specified primitives. This latter feature distinguishes the approach from other so-called neuro-symbolic approaches (see Box 2 for further discussion). In the following sections, we first provide a general characterization of this approach, drawing on concepts from information theory, and discuss a number of recently proposed neural network architectures that implement the approach. We then discuss the potential of the approach for modeling human cognition, relating it to existing cognitive theories and considering potential mechanisms through which it might be implemented in the brain.
@@
We define the relational bottleneck as any mechanism that restricts the flow of information from perceptual to downstream reasoning systems to consist only of relations (see Box 1 for a formal definition). For example, consider the relational patterns depicted in Figure 1. In this example, the images on the left and the images on the right are governed by the same abstract relational pattern (ABA), but these patterns are instantiated by different sets of objects (dogs and cats on the left, planes and cars on the right). Given inputs representing the individual objects (e.g., representations of the dog and cat images on the left), a relational bottleneck would constrain the representations passed to downstream reasoning processes such that they capture only the specific relations between these objects (e.g., whether the objects have the same shape), and discard information about the features of the individual objects (e.g., information about dogs and cats). The result is that the images on the left will have the same representation as the images on the right, despite the different objects depicted in these images. Such a representation encourages downstream processes to identify relational patterns in a manner that is abstracted away from the details of specific instances of those patterns, and can therefore be systematically generalized to novel inputs (see Concluding Remarks for discussion of non-relational factors in human cognition, and how these might be integrated with the relational bottleneck). In the following section, we highlight three recently proposed neural network architectures that instantiate this approach in different guises, illustrating how they utilize a relational bottleneck to induce abstract concepts in a data-efficient manner.
Box 1: The relational bottleneck principle
@@
where $r(x_{i},x_{j})$ is a learned relation function that satisfies certain key relational properties (e.g., transitivity). One type of operation that satisfies the relevant properties is inner products of the form $\langle\phi(x_{i}),\phi(x_{j})\rangle$. Let $\mathcal{R}$ be the class of all possible relational representations of the input signal $X$. In a relational task, there exists $R\in\mathcal{R}$ which is sufficient for predicting $Y$ (i.e., $X\to R\to Y$).
A relational bottleneck is any mechanism that restricts the space of all possible compressed representations to be a subset of the relational signals $\mathcal{R}$. This gives the model a smaller space of possible compressed representations over which it must search. This space of compressed representations $\mathcal{R}$ is guaranteed to contain a sufficient representation for the task and excludes many representations that encode extraneous information about $X$, promoting efficient learning of relational abstractions.
@@
Figure 2 (Key Figure) depicts three neural architectures that implement the relational bottleneck through the use of architectural inductive biases. Here, we discuss how the distinct mechanisms employed by these models implement the same underlying principle. In particular, a common aspect of all three architectures is the use of inner products to represent relations, which ensures that the resulting representations are genuinely relational. In each case, we also contrast these architectures with closely related approaches that do not incorporate a relational bottleneck, emphasizing how this key architectural feature gives rise to the data-efficient induction of abstractions.
@@
We first consider the Emergent Symbol Binding Network (ESBN) (Figure 2[34]. The ESBN is a deep neural network architecture, augmented by an external memory (see Glossary for general definition, details of architecture explained below), that was inspired by the notion of role-filler variable binding in cognitive models of relational reasoning [37, 38, 39]. In those models, relational reasoning is supported by the existence of separately represented ‘roles’, which represent information about abstract variables, and ‘fillers’, which capture information about concrete entities to which those variables are bound. This coding scheme enables roles and fillers to be flexibly combined in new ways, thus capturing one of the key properties of symbol-processing systems: the ability of symbolic variables to be associated with any potential values. Previous work has focused on how the binding might be performed in neural circuits (e.g., units that code for a particular role may be temporarily ‘bound’ to units that code for a particular filler by firing in synchrony with one another [38]). However, the role and filler representations themselves were typically pre-specified by the modeler, leaving open the question of how such symbolic representations might emerge from experience.
@@
The ESBN adopts this key idea of separate role and filler representations, but integrates them into a larger system that can be trained end-to-end (via backpropagation), averting the need to pre-specify either type of representation. The ESBN contains three major components: 1) a feedforward encoding pathway (‘Encoder’ in Figure 2), which generates object embeddings (corresponding to fillers) from high-dimensional perceptual inputs (e.g., images); 2) a recurrent controller (‘Controller’ in Figure 2), which operates over learned representations of abstract task-relevant variables (corresponding to roles; though the system is not explicitly trained to represent any particular task-relevant variables, but instead learns these representations through backpropagation); and 3) an external memory system responsible for binding and associating representations of roles and fillers (i.e., abstract variables and perceptual embeddings) between the two pathways. The ESBN processes pairs of representations (perceptual observations and abstract variables) sequentially. For each observation, a pair of representations is added to memory (by appending them to the list of memories from previous timesteps), one from the perceptual pathway (referred to as a key), and one from the control pathway (referred to as a value; note that the use of the terms ‘key’ and ‘value’ here is reversed relative to the original paper [34] in order to be more consistent with their usage in describing the CoRelNet and Abstractor architectures). To read from this memory, the object embedding for the current observation (referred to as a query) is compared to all of the keys in memory via an inner product, yielding a set of scores (one for each key) that indicate the similarity of the current object embedding (i.e., filler) to each entry in memory. These scores are then used to compute a weighted average of the value embeddings (i.e., roles) in the abstract pathway, which is then retrieved and passed to the controller.
[…]
@@
Figure 3: Results. (a) Results for the ESBN and baseline architectures (Transformer, Neural Turing Machine (NTM), Metalearned Neural Memory (MNM), Long Short-Term Memory (LSTM), PrediNet, and the Relation Network (RN)) on the identity rules task, reproduced from [34]. X axis represents the number of potential objects (out of 100 possible objects) withheld during training. When all objects are observed during training (0 withheld), most baselines perform well on the task. When most objects are withheld (95 withheld; test set includes only objects withheld during training), only the ESBN generalizes well to new objects. (b) Results for an object sorting task involving an asymmetric relation (greater-than/less-than), reproduced from [36]. The abstractor learns this task significantly faster than both the transformer and an ablation model in which relational cross-attention is replaced by standard cross-attention. (c) Results from the give-N task, reproduced from [40]. X axis represents the target number N (desired number of objects). Y axis represents the episode at which the model reaches a particular criterion for the ability to count to each value of N. ESBN learns the task significantly faster than the LSTM or Transformer baselines. ESBN also displays inductive transition (rapid learning for $N>5$) similar to that observed in human development.
@@
Importantly, in this retrieval operation, the control pathway cannot access the content of the representations in the perceptual pathway (i.e., the fillers). Instead, the interaction is mediated only by the comparison of perceptual representations with each other. In other words, to the extent that a particular perceptual memory embedding (key) is similar to the current perceptual embedding (filler, used as a query), the corresponding control memory (value) will be retrieved and passed to the controller to be used as the role; critically, the perceptual embedding itself will not be passed to the controller. The ESBN thus implements the relational bottleneck as an architectural prior, separating the learning and use of abstract representations (roles) by the controller from the embeddings of perceptual information (fillers). It is precisely this separation that guarantees that the representations in the control pathway are abstract. Thanks to this design feature, the ESBN is capable of rapidly learning relational patterns (such as the identity rules displayed in Figure 1), and generalizing them to out-of-distribution inputs (e.g., to previously unseen shapes; Figure 3[34]. Critically, it can be shown to use precisely the same representation for a given role, irrespective of filler, thus exhibiting a critical feature of abstract, symbolic processing [13]. In this sense, the representations in the model’s control pathway can be viewed as a form of learned ‘symbols’.
@@
It is instructive to compare this model with similar approaches that do not implement the relational bottleneck. The ESBN is part of a broader family of neural network architectures that use content-addressable external memory – a separate store of information with which a neural network can interact via learnable read and write operations [41, 42, 43]. Notably, these read and write operations typically rely on a similarity computation (based on inner products). These have often been cast as simplified implementations of the brain’s episodic memory system [44, 45]. Standard external memory architectures do not typically isolate the control and perceptual pathways. Instead, perceptual inputs are passed directly to a central controller, which is then responsible for writing to and reading from a single, monolithic memory. Though it is possible for a role-filler structure to emerge in these systems given a sufficiently large amount of training data [46], they take much longer to learn relational tasks (requiring approximately an order of magnitude more training data), and do not display the same degree of generalization as the ESBN [34]. Thus, although external memory plays an important role in the ESBN, the presence of external memory alone is insufficient to enforce a relational bottleneck. Rather, it is the isolation of the perceptual and abstract processing components from one another that does so. Furthermore, as we illustrate in the following sections, it is possible to achieve this isolation without the use of external memory.
@@
An alternative approach to implementing the relational bottleneck is illustrated by the Compositional Relation Network (CoRelNet) (Figure 2[35]. In that approach, a set of perceptual observations are first processed by an encoder, yielding a sequence of object embeddings. A relation matrix is then computed over all pairs of objects (within a given instance of a problem, e.g., all pairs of objects found within the ABA pattern in Figure 1), in which each entry consists of the inner product between a pair of object embeddings (thus capturing the similarity between each pair of objects). Finally, this relation matrix is passed to a downstream decoder network (the architecture of this network can vary, e.g., using a multilayer perceptron or transformer). The relation matrix thus forms a relational bottleneck: all perceptual information flows into this matrix, converting it into a form that only preserves relational information (about the similarity between objects), and only this relational information is then passed on to the decoder. As with the ESBN, this relational bottleneck enables CoRelNet to rapidly learn and systematically generalize relational patterns.
CoRelNet can be viewed as a feedforward, parallelized implementation of the sequential process (of encoding and similarity-based retrieval from external memory) carried out by the ESBN. This results in performance benefits, as CoRelNet does not suffer from the vanishing gradient problem that is a challenge for recurrent neural networks used to implement sequential processing [47] (due to the fact that relations are processed in parallel, rather than sequentially as in the ESBN). It also makes the key relational inductive bias underlying the ESBN more explicit. The ESBN’s memory retrieval procedure, in which the current observation is compared to the entries in memory, can be interpreted as computing a single row of the relation matrix. In both architectures, downstream processing is constrained so as to depend only on this relation matrix (which forms the relational bottleneck), though the details of this dependency differ.
@@
The recently proposed Abstractor architecture (Figure 2[36] illustrates how the relational bottleneck can be implemented within the broader framework of attention-based architectures (including the Transformer [52]). The Abstractor is built on a novel attention operation termed relational cross-attention. In this operation, a set of object embeddings (which may be produced by an encoder given perceptual observations) is converted to form keys and queries, using separate linear projections. A relation matrix is then computed, in which each entry corresponds to the inner product between a query and key. The relation matrix is used to attend over a set of learned values, a separate set of representations that are learned through backpropagation (based on the downstream task for which the network is trained) which reference objects but are independent of their attributes. In this attention operation, each value representation is replaced with a weighted average of the set of all value representations, where the weights are determined by the entries of the relation matrix. As with the CoRelNet architecture, the relation matrix thus forms a relational bottleneck, in the sense that it converts all perceptual information to relational information, and downstream processing (in this case, the relational cross-attention operation) then depends only on this relational information.
@@
This implementation of the relational bottleneck yields the same benefits observed in others: the Abstractor learns relational patterns faster than the Transformer, and displays better out-of-distribution generalization of those patterns. The Abstractor also has a few advantages relative to existing implementations of the ESBN and CoRelNet. Because the relation matrix is computed using separate key and query projections, the Abstractor is capable of representing asymmetric relations (e.g., can capture the difference in meanings between ‘A is greater than B’ and ‘B is greater than A’; Figure 3). In addition, multi-headed relational cross-attention enables the Abstractor to model multi-dimensional relations. As proposed, ESBN and CoRelNet are limited to relations along a single feature dimension only. Finally, similar to Transformers, the Abstractor is a generative architecture, whereas the ESBN and CoRelNet are purely discriminative (although an alternative implementation of the ESBN has been proposed that can perform generative tasks [53]). This enables the Abstractor to perform a broader range of tasks, including the sequence-to-sequence tasks that are common in natural language processing.
@@
The neural network modeling work discussed in the previous sections was focused primarily, though not exclusively, on relational patterns involving same/different relations. Although similarity is fundamental to human reasoning, and has indeed been a central focus of theories of relational and analogical reasoning [54, 55, 56], human cognition is also characterized by more complex relation types, including asymmetric relations [57] and higher-order relations [5] (relations between relations). It is worth emphasizing that the relational bottleneck can also account for these more complex relation types. First, although the inner product between two vectors is inherently symmetric ($\mathbf{a}\cdot\mathbf{b}$ is identical to $\mathbf{b}\cdot\mathbf{a}$), the relational bottleneck can also account for asymmetric relations by computing the inner product between separate key and query embeddings ($\mathbf{q_{a}}\cdot\mathbf{k_{b}}$ is not the same as $\mathbf{q_{b}}\cdot\mathbf{k_{a}}$). This use of separate key and query embeddings enables the Abstractor to model asymmetric relations such as greater-than/less-than, as well as other asymmetric relations found in mathematical reasoning problems [36] (and a similar architectural modification is also possible for the ESBN and CoRelNet). Second, although the architectures discussed here focused on pairwise relations, higher-order relations can be straightforwardly accommodated through the recursive application of the relational bottleneck (i.e., by treating the outputs of one relational bottleneck as the inputs to another relational bottleneck, thus computing relations between other relations). Along these lines, recent work proposed relational convolutional networks [58] and demonstrated that a compositional relational architecture can learn representations of higher-order relations, outperforming both non-compositional relational architectures (e.g., CoRelNet) and deep non-relational architectures (e.g., transformers). This suggests that relational inductive biases and compositionality are both necessary for efficiently learning representations of higher-order relations.

### v5

@@ Abstract
A central challenge for cognitive science is to explain how abstract concepts are acquired from limited experience. This has often been framed in terms of a dichotomy between connectionist and symbolic cognitive models. Here, we highlight a recently emerging line of work that suggests a novel reconciliation of these approaches, by exploiting an inductive bias that we term the relational bottleneck. In that approach, neural networks are constrained via their architecture to focus on relations between perceptual inputs, rather than the attributes of individual inputs. We review a family of models that employ this approach to induce abstractions in a data-efficient manner, emphasizing their potential as candidate models for the acquisition of abstract concepts in the human mind and brain.
@@
  • Human learners acquire abstract concepts from limited experience. The effort to explain this capacity has fueled debate between symbolic and connectionist approaches, and motivated proposals for neuro-symbolic systems.
  • The relational bottleneck principle suggests a novel way to bridge the gap. By restricting information processing to focus only on relations, the approach encourages abstract symbol-like mechanisms to emerge in neural networks.
  • We present an information theoretic formulation, and review neural network architectures that implement the principle, enabling rapid learning and systematic generalization of relational patterns.
  • The approach can explain phenomena ranging from the development of numerical abstractions to capacity limits in cognition; is consistent with findings from cognitive neuroscience; and offers a principle for designing more powerful artificial learning systems.
@@
Human cognition displays a remarkable ability to transcend the specifics of limited experience to entertain highly general, abstract ideas. Understanding how the mind and brain accomplish this has been a central challenge of cognitive science, and a major preoccupation of philosophy before that [1, 2, 3, 4]. Of particular importance is the central role played by relations, which enable human reasoners to abstract away from individual entities and identify higher-order patterns across distinct domains [5, 6]. For instance, when presented with the images in Figure 1, one can easily determine that a common relational pattern (ABA) is displayed on both the left and right, despite the involvement of completely different objects. This capacity is a major component underlying the human capacity for fluid reasoning [7, 8], and has been proposed as a key factor distinguishing human intelligence from that of other species [9].
@@
A long tradition in cognitive science and AI [10, 11, 12] holds that this capacity for abstraction depends on processes akin to symbolic programs. A major appeal of this approach is that symbols are, by design, abstracted away from the content to which they refer, thus naturally accounting for the flexibility and systematicity of human concepts [13]. More recently, program induction (see Glossary) has provided an account of how symbolic concepts might be learned directly from data [14, 15, 16, 17, 18, 19, 20], formalizing learning as a search for the program that maximizes the likelihood of observed data. However, while this approach is capable in principle of representing any possible set of concepts [21], the discovery of these concepts using traditional search methods often proves intractable, making it difficult in practice to identify programs with the richness and complexity of human natural concepts (though see Box 1 for discussion).
@@
An alternative approach, connectionism, has for decades explored how cognitive abstractions might emerge through experience in general-purpose neural architectures [22, 23, 24, 25]. This endeavor has taken on new relevance with the advent of large language models, demonstrating that it is possible, in some cases, for a human-like capacity for abstraction to emerge given sufficient scaling of architecture and training data [26, 27, 28, 29]. For instance, it has recently been shown that large language models can solve various analogy problems at a level equal to that of college students [30]. However, the ability of these models to perform abstract tasks depends on exposure to a much larger training corpus than individual humans receive in an entire lifetime [31, 32], thus failing to account for the data efficiency of human concept learning.
@@
In this review, we highlight an emerging approach that suggests a novel reconciliation of these two traditions. The central feature of this approach is an inductive bias that we refer to as the relational bottleneck: a constraint that biases neural network models to focus on relations between objects rather than the attributes of individual objects. This approach enables the data efficiency associated with symbolic cognitive models, while retaining the scalable training procedures associated with neural network models (see Box 1 for further discussion of neuro-symbolic approaches). In the following sections, we first provide a general characterization of this approach, drawing on information theory, and discuss recently proposed neural network architectures that implement the approach. We then discuss the potential of the approach for modeling human cognition, relating it to existing theories and considering mechanisms through which it might be implemented in the brain.
Box 1: Neuro-symbolic modeling approaches
More recently, hybrid systems have been developed that combine deep learning with symbolic programs. In this approach, deep learning components are typically employed to translate raw perceptual inputs, such as images or natural language, into symbolic representations, which can then be processed by traditional symbolic algorithms [40, 41, 42]. This approach is complemented by recent neuro-symbolic approaches to probabilistic program induction, in which symbolic primitives are pre-specified (following earlier symbolic-connectionist modeling efforts), and then deep learning is used to assemble these primitives into programs [17].
@@
We define the relational bottleneck as any mechanism that restricts the flow of information from perceptual to downstream reasoning systems to consist only of relations (see Box 2 for a formal definition). For example, consider the relational patterns depicted in Figure 1. In this example, the images on the left and right are governed by the same abstract pattern (ABA), but these patterns contain different sets of objects (dogs and cats vs. planes and cars). Given inputs representing individual objects (e.g., representations of the dog and cat images on the left), a relational bottleneck would constrain the representations passed to downstream reasoning processes to capture only the relations between these objects (e.g., whether the objects have the same shape), and discard information about the individual objects (e.g., information about dogs and cats). The result is that the images on the left will have the same representation as those on the right, despite the different objects depicted in these images. This encourages downstream processes to identify relational patterns in a manner that is abstracted away from specific instances of those patterns, and can therefore be systematically generalized to novel inputs (see Concluding Remarks for discussion of non-relational factors in human cognition). In the following section, we highlight three recently proposed neural architectures that instantiate this approach in different guises, illustrating how they utilize a relational bottleneck to induce abstract concepts in a data-efficient manner. It is also worth emphasizing that, although we focus here on artificial neural networks, the approach may also be applicable to other modeling approaches (see Outstanding Questions).
Box 2: The relational bottleneck principle
@@
where $r(x_{i},x_{j})$ is a relation function. A “relation function” is a function that takes a pair of objects as input and returns a relation between them (see ‘Modeling more complex relations’ for discussion of higher-order relations). We say that a task is relational if there exists some relational signal $R$ which is sufficient for predicting the target $Y$ (i.e., $X\to R\to Y$). A relational bottleneck is any mechanism that restricts or biases the learned compressed representation of the input to be a relational representation, separated from object-level features. This gives the model a smaller space of possible compressed representations over which it must search. Moreover, this restricted space is guaranteed to contain a sufficient representation for the task and excludes many representations that encode extraneous information about $X$, promoting efficient learning of relational abstractions.
[…]
@@
Figure 2 (Key Figure) depicts three architectures that implement the relational bottleneck through architectural inductive biases. Here, we discuss how the distinct mechanisms in these models implement the same underlying principle. In particular, a common element is the use of inner products to represent relations, which ensures that the resulting representations are genuinely relational. In each case, we also contrast these architectures with related approaches that do not incorporate a relational bottleneck, emphasizing how this key architectural feature enables the data-efficient induction of abstractions.
@@
We first consider the Emergent Symbol Binding Network (ESBN) (Figure 2(a)[52], a deep neural network architecture inspired by the notion of role-filler variable binding in cognitive models of relational reasoning [34, 33, 55]. In those models, relational reasoning is supported by the existence of separate ‘roles’, representing information about abstract variables, and ‘fillers’, representing information about concrete entities bound to those variables. This coding scheme enables roles and fillers to be flexibly combined in new ways, capturing a key property of symbol-processing systems: the ability of symbolic variables to be associated with any potential values. Previous work focused on how binding might be performed in neural circuits (e.g., units coding for a particular role may be temporarily ‘bound’ to units coding for a particular filler by firing synchronously [33]). However, role and filler representations were typically pre-specified by the modeler, leaving open the question of how such symbolic representations might be learned.
@@
The ESBN adopts this key idea of separate roles and fillers, but integrates them into a system that can be trained end-to-end (via backpropagation), averting the need to pre-specify these representations. The ESBN contains three major components: 1) a feedforward encoding pathway (‘Encoder’ in Figure 2(a)), which generates object embeddings (i.e., fillers) from perceptual inputs (e.g., images); 2) a recurrent controller (‘Controller’ in Figure 2(a)), which operates over learned representations of abstract variables (i.e., roles; though the system is not explicitly trained to represent any particular variables, but instead learns these representations through backpropagation); and 3) an external memory system responsible for binding representations of roles and fillers (i.e., abstract variables and perceptual embeddings). The ESBN processes inputs sequentially. For each observation, a pair of representations is appended to memory, one from the perceptual pathway (referred to as a key), and one from the control pathway (referred to as a value; note that the use of the terms ‘key’ and ‘value’ here is reversed relative to the original paper [52] in order to be consistent with their usage in describing the CoRelNet and Abstractor architectures). To read from this memory, the embedding for the current observation (referred to as a query) is compared to all keys in memory via an inner product, yielding a set of scores (one for each key) that indicate the similarity of the current object embedding (i.e., filler) to each entry in memory. These scores are then used to compute a weighted average of the value embeddings (i.e., roles) in the abstract pathway, which is then retrieved and passed to the controller.
@@
Figure 3: The relational bottleneck encourages data-efficient and generalizable relation learning. (a) Results for the ESBN and baseline architectures (Transformer, Neural Turing Machine (NTM), Metalearned Neural Memory (MNM), Long Short-Term Memory (LSTM), PrediNet, and the Relation Network (RN)) on the identity rules task, reproduced from [52]. X axis represents the number of potential objects (out of 100 possible objects) withheld during training. When all objects are observed during training (0 withheld), most baselines perform well on the task. When most objects are withheld (95 withheld; test set includes only objects withheld during training), only the ESBN generalizes well to new objects. (b) Results for an object sorting task involving an asymmetric relation (greater-than/less-than), reproduced from [54]. The abstractor learns this task significantly faster than both the transformer and an ablation model in which relational cross-attention is replaced by standard cross-attention. (c) Results from the give-N task, reproduced from [56]. X axis represents the target number N (desired number of objects). Y axis represents the episode at which the model reaches a particular criterion for the ability to count to each value of N. ESBN learns the task significantly faster than the LSTM or Transformer baselines. ESBN also displays inductive transition (rapid learning for $N>5$) similar to that observed in human development.
@@
Importantly, in this retrieval operation, the control pathway cannot access the content of the representations in the perceptual pathway (i.e., the fillers). Instead, the interaction is mediated only by the comparison of perceptual representations with each other. In other words, to the extent that a particular perceptual memory embedding (key) is similar to the current perceptual embedding (filler, used as a query), the corresponding control memory (value) will be retrieved and passed to the controller to be used as the role; critically, the perceptual embedding itself is not passed to the controller. The ESBN thus implements the relational bottleneck as an architectural prior, separating the learning and use of abstract representations (roles) by the controller from the embeddings of perceptual information (fillers). It is precisely this separation that guarantees the representations in the control pathway are abstract. This feature enables the ESBN to rapidly learn relational patterns (such as the identity rules displayed in Figure 1), and generalize them to out-of-distribution inputs (e.g., previously unseen shapes; Figure 3(a)[52]. Indeed, the ESBN can learn relational patterns from only a handful of examples (as few as 4 examples), mirroring the data efficiency of relational learning in young children (who can typically learn to perform relational tasks with fewer than 20$\sim$ 20∼ 20 examples [57], as opposed to the thousands of examples that are necessary for standard neural network architectures). Critically, the ESBN can be shown to use precisely the same representation for a given role, irrespective of filler, thus exhibiting a critical feature of abstract, symbolic processing [13]. In this sense, the representations in the model’s control pathway can be viewed as a form of learned ‘symbols’.
@@
It is instructive to compare this model to similar approaches without a relational bottleneck. The ESBN is part of a broader family of architectures that use content-addressable external memory – a separate store of information with which a neural network can interact via learnable read and write operations [58, 59]. Notably, these read and write operations rely on a similarity computation (based on inner products). These have often been cast as simplified implementations of the brain’s episodic memory system [60, 61]. Standard external memory architectures do not typically isolate the control and perceptual pathways. Instead, perceptual inputs are passed directly to a central controller, which is responsible for writing to and reading from a single, monolithic memory. Though it is possible for a role-filler structure to emerge in these systems given a sufficiently large amount of training data [62], they take much longer to learn relational tasks (requiring approximately an order of magnitude more training data), and do not generalize as well as the ESBN [52]. Thus, although external memory plays an important role in the ESBN, the presence of external memory alone is insufficient to enforce a relational bottleneck. Rather, it is the isolation of the perceptual and abstract processing components from one another that does so. Furthermore, as we illustrate in the following sections, it is possible to achieve this isolation without the use of external memory.
@@
An alternative approach to implementing the relational bottleneck is illustrated by the Compositional Relation Network (CoRelNet) (Figure 2(b)[53]. In that approach, a set of perceptual observations are first processed by an encoder, yielding a sequence of object embeddings. A relation matrix is then computed over all pairs of objects (within a given problem instance, e.g., all pairs of objects found within the ABA pattern in Figure 1), in which each entry consists of the inner product between a pair of object embeddings (thus capturing the similarity between each pair of objects). Finally, this relation matrix is passed to a downstream decoder network (the architecture of this network can vary, e.g., using a multilayer perceptron or transformer). The relation matrix thus forms a relational bottleneck: all perceptual information flows into this matrix, converting it into a form that only preserves relational information (about the similarity between objects), and only this relational information is then passed on to the decoder. As with the ESBN, this relational bottleneck enables CoRelNet to rapidly learn and systematically generalize relational patterns.
CoRelNet can be viewed as a feedforward, parallelized implementation of the sequential process (of encoding and similarity-based retrieval from external memory) carried out by the ESBN. This results in performance benefits, as CoRelNet does not suffer from the vanishing gradient problem that is a challenge for recurrent neural networks used to implement sequential processing [63]. That is, the gradients associated with all relations in a given scene are processed in parallel via a single step of backpropagation, rather than being diluted over several iterations, as is the case for the sequential processing carried out by the ESBN. This approach also makes the key relational inductive bias underlying the ESBN more explicit. The ESBN’s memory retrieval procedure, in which the current observation is compared to the entries in memory, can be interpreted as computing a single row of the relation matrix. In both architectures, downstream processing is constrained so as to depend only on this relation matrix (which forms the relational bottleneck), though the details of this dependency differ.
@@
The recently proposed Abstractor architecture (Figure 2(c)[54] illustrates how the relational bottleneck can be implemented within the broader framework of attention-based architectures (including the Transformer [67]). The Abstractor is built on a novel attention operation termed relational cross-attention. In this operation, a set of object embeddings (which may be produced by an encoder given perceptual observations) is converted to form keys and queries, using separate linear projections. A relation matrix is then computed, in which each entry corresponds to the inner product between a query and key. The relation matrix is used to attend over a set of learned values. The attention operation itself is identical to standard self-attention, in which each value embedding is replaced with a weighted average of the set of all value embeddings, where the weights are determined by the match between queries and keys (i.e., the relation matrix). However, the value embeddings used for relational cross-attention are formed from a separate set of representations that are learned through backpropagation (based on the downstream task for which the network is trained) which reference objects but are independent of their attributes (in the simplest scheme, these values reference objects via their position, though more sophisticated referencing schemes are also possible  [54]). This is in contrast to standard self-attention, in which the values are computed via a linear projection of the perceptual inputs (just as the queries and keys are computed). As with the CoRelNet architecture, the relation matrix thus forms a relational bottleneck, in the sense that it converts all perceptual information to relational information, and downstream processing (in this case, the relational cross-attention operation) then depends only on this relational information.
@@
This implementation of the relational bottleneck yields the same benefits observed in others: the Abstractor learns relational patterns faster than the Transformer, and displays better out-of-distribution generalization of those patterns. The Abstractor also has a few advantages relative to existing implementations of the ESBN and CoRelNet. Because the relation matrix is computed using separate key and query projections, the Abstractor is capable of representing asymmetric relations (e.g., can capture the difference in meanings between ‘A is greater than B’ and ‘B is greater than A’; Figure 3(b)). In addition, multi-head relational cross-attention enables the Abstractor to model multi-dimensional relations. As proposed, ESBN and CoRelNet are limited to relations along a single feature dimension only. Finally, similar to Transformers, the Abstractor is a generative architecture, whereas the ESBN and CoRelNet are purely discriminative (although an alternative implementation of the ESBN has been proposed that can perform generative tasks [68]). This enables the Abstractor to perform a broader range of tasks, including the sequence-to-sequence tasks that are common in natural language processing.
@@
The neural network modeling work discussed in the previous sections was focused primarily, though not exclusively, on relational patterns involving same/different relations. Although similarity is fundamental to human reasoning, and has indeed been a central focus of theories of relational and analogical reasoning [69, 70, 71], human cognition is also characterized by more complex relation types, including asymmetric relations [72] and higher-order relations [5] (relations between relations). It is worth emphasizing that the relational bottleneck can also account for many of these more complex relation types. First, although the inner product between two vectors is inherently symmetric ($\mathbf{a}\cdot\mathbf{b}$ is identical to $\mathbf{b}\cdot\mathbf{a}$), the relational bottleneck can also account for asymmetric relations by computing the inner product between separate key and query embeddings ($\mathbf{q_{a}}\cdot\mathbf{k_{b}}$ is not the same as $\mathbf{q_{b}}\cdot\mathbf{k_{a}}$). This use of separate key and query embeddings enables the Abstractor to model asymmetric relations such as greater-than/less-than, as well as other asymmetric relations found in mathematical reasoning problems [54] (and a similar architectural modification is also possible for the ESBN and CoRelNet). Second, although the architectures discussed here focused on pairwise relations, higher-order relations can be straightforwardly accommodated through the recursive application of the relational bottleneck (i.e., by treating the outputs of one relational bottleneck as the inputs to another relational bottleneck, thus computing relations between other relations). Along these lines, recent work proposed relational convolutional networks [73] and demonstrated that a hierarchical relational architecture can learn representations of higher-order relations, outperforming both non-hierarchical relational architectures (e.g., CoRelNet) and deep non-relational architectures (e.g., transformers). These results illustrate how the relational bottleneck can account for more complex relation types, but it remains an important avenue for future work to investigate whether and how such architectures can account for the full space of relations that characterize human cognition.